OpenStudy (abbyrose0105):

For the principal quantum number n=3, how many possible values are there for the angular momentum quantum number (l)?

OpenStudy (abbyrose0105):

My options are: 2 3 6 8

OpenStudy (aaronq):

\(l=n-1>0\) integers only.

OpenStudy (aaronq):

\(l=3-1=2,~1,~0\)

OpenStudy (abbyrose0105):

So the answer is 2?

OpenStudy (abbyrose0105):

@aaronq

OpenStudy (aaronq):

those are values, there are 3 possible values, 2, 1 and 0.

OpenStudy (abbyrose0105):

Oh ok got it